Which pickup should be purchased by Northwest Construction for use by its job superintendent? How much does this choice save (using either PW or EAC)? Northwest uses a 12% rate to evaluate its capital investments. What assumptions have you made, and why?

Pickup First Cost Maintenance Cost/Year Salvage Value Life
New $18000 $300 $5000 5
Old 10000 750 3000 3
Answer #1

EAC of new pickup = 18000 * (A/P, 12%,5) + 300 - 5000 * (A/F, 12%,5)

= 18000 * 0.277410 + 300 - 5000 *0.157410

= 4506.33

EAC of old pickup = 10000 * (A/P, 12%,3) + 750 - 3000 * (A/F, 12%,3)

= 10000 *0.416349 + 750 - 3000 *0.296349

= 4024.44

As the EAC of old pickup is less, it should be selected